Sell2All, Inc. accounts for its revenues under the installment sales method. In year​ 1, Sell2All sold inventory with a cost of​ $300,000 for​ $400,000 and collected​ $100,000. In year​ 2, Sell2All sold inventory with a cost of​ $500,000 for​ $750,000 and collected​ $400,000, including​ $100,000 related to the year 1 sales and​ $300,000 related to the year 2 sales. What amount of earned gross profit should Sell2All report on its December​ 31, year 2 balance sheet?

A.$400,000

B.​$100,000

C.​$125,000

D.​$133,200




1
Expert's answer
2021-04-29T10:30:41-0400

Solution:

The correct answer is C. $125,000

Sell2All Inc. should report earned gross profit of $125,000 as at December 31st, year 2 balance sheet.

 

Calculate gross profit margin:


Gross profit margin="\frac{Revenue - Cost\; of\; sales}{Revenue} \times 100"


Year 1 Gross profit = "\frac{400,000 - 300,000}{400,000}\times 100 = 25\%"

Year 1 earned gross profit = Amount collected "\times" gross profit margin

= 100,000"\times" "25\%" = "\$25,000"

Year 2 earned gross profit = 400,000 "\times" "25\%" = "\$100,000"


Total gross profit earned to be recorded on the year 2 balance sheet as at December 31st = 25,000 + 100,000 = "\$125,000"